72 If f(a) t dt then 3 W f(x) = Check Answer
Answer #1

Leibniz Inegral Rule х X h) d da f(t) at gia) da - (+) ) figaro) de g(0) dx

So, for 25 3 f1u); t at -3 f(X) = d.f(x) da

2 =d 3 dt da -3 = (x²3 d d (x²) -30°d (3) dx 3 dx 6 = x (2x) -(-3)*(0) = 2 xc7

7 .: (c)=127 x)